Please solve accordingly.

Problem 12) Relation between Angular and Linear Quantities II A centrifuge can produce a radial acceleration of 3000g when rotating at an angular speed of 5000 rev/min. Calculate the required radius of the centrifuge.
Answer #1

We know that,

v = rw

a = v^2/r = mg

mg = (rw)^2/r

r = mg/w^2

where, w = (5000rev/min)*(2πrad/rev)*(1min/60s)

= 5000*2*pi/60 = 523.6 rad/s

r = 3000 * 9.8/(523.6)^2

= 0.107 m